cooling: 1) cover - hoses to distribution: C-side done, incl leak test 2) inner heat screen - installation of distribution, hoses, and leak test 1.5 t 3) pressure relief valve - install, with plug for pressure sensors, hoses: done, except near A00,A05 4) leak and pressure test (up to 300 mbar) for all sectors: next week 5) NPT adapters for pressure relief valves: 14.12. delivery. Pressure gauges end Jan 6) ss adapters for all cooling connections: offer not yet received 6) leak hunting: 2x chamber cooling circuits still not ok (A, C) FEC: 1) change FEC A05, check electronics: done 2) A13, check electronics: done 3) Wiener test: next week cabling: 1) C12 gating test: open (Danilo) 2) connect patch panel cables (cover, edge) Rainer, Joachim, A-side done 3) redo skirt cable routing and testing: open, part in cavern 07 4) storage and fixation of cable rolls: Rainer, technicians 5) CAN-ELMB cable: replace with shielded: Rainer --> 07 6) ELMB adapter resistor: Joachim, Marco, done?? 7) connector at bretelle display: Thomas: next week misc: 1) ITS-rails: done 2) ITS cameras: 18.12., Bjorn 3) orientation of bretelles (Spanner)?